# Quadratic Gating Functions in Mixture of Experts: A Statistical Insight

### v1

@@ Appendix D Proof of Theoretical Results
which indicates that $\|f_{G}-f_{\check{G}}\|_{L_{2}(\mu)}\leq L_{1}k\tau$. Here, the second inequality is according to the Cauchy-Schwarz inequality, the third inequality occurs as the softmax weight is bounded by 1, and the fourth inequality follows from the fact that the expert $h({\bm{x}},\cdot)$ is a Lipschitz function with Lipschitz constant $L_{1}$. Next, we have
$\|f_{\check{G}}-f_{\overline{G}}\|_{L_{2}(\mu)}^{2}$ $\leq k\int\sum_{i=1}^{k}\Bigg{[}\Bigg{(}\frac{\exp({\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+({\bm{b}}_{i})^{\top}{\bm{x}}+c_{i})}{\sum_{j=1}^{k}\exp({\bm{x}}^{\top}A_{j}{\bm{x}}+(b_{j})^{\top}{\bm{x}}+c_{j})}$
$\hskip 71.13188pt-\frac{\exp({\bm{x}}^{\top}\overline{A}_{i}{\bm{x}}+(\overline{b}_{i})^{\top}{\bm{x}}+\overline{c}_{i})}{\sum_{j=1}^{k}\exp({\bm{x}}^{\top}\overline{A}_{j}{\bm{x}}+(\overline{b}_{j})^{\top}{\bm{x}}+\overline{c}_{j})}\Bigg{)}\cdot h({\bm{x}},\overline{{\bm{\eta}}}_{i})\Bigg{]}^{2}~{}\mathrm{d}\mu({\bm{x}})$
$\leq kM^{2}L^{2}\int\sum_{i=1}^{k}\Big{[}\|A_{i}-\overline{A}_{i}\|\cdot\|x\|^{2}+\|{\bm{b}}_{i}-\overline{b}_{i}\|\cdot\|x\|+|c_{i}-\overline{c}_{i}|\Big{]}^{2}\mathrm{d}\mu({\bm{x}})$
$\leq kM^{2}L^{2}\int\sum_{i=1}^{k}(\tau B^{2}+\tau B+\tau)^{2}\mathrm{d}\mu({\bm{x}})$
$\leq[kML\tau(B^{2}+B+1)]^{2},$
where $L\geq 0$ is a Lipschitz constant of the softmax weight. This result implies that $\|f_{\check{G}}-f_{\overline{G}}\|_{L_{2}(\mu)}\leq kML(B^{2}+B+1)\tau$. According to the triangle inequality, we have
$\|f_{G}-f_{\overline{G}}\|_{L_{2}(\mu)}\leq\|f_{G}-f_{\check{G}}\|_{L_{2}(\mu)}+\|f_{\check{G}}-f_{\overline{G}}\|_{L_{2}(\mu)}\leq[L_{1}k+kML(B^{2}+B+1)]\cdot\tau.$
@@
${N}(\tau,\mathcal{F}_{k}(\Theta),\|\cdot\|_{L_{2}(\mu)})\leq|\Delta_{\tau}|\times|\Omega_{\tau}|\leq\mathcal{O}_{P}(n^{-(d^{2}+d+1)k})\times\mathcal{O}(n^{-qk})\leq\mathcal{O}(n^{-(d^{2}+d+1+q)k}).$ (31)
@@
$H_{B}(4\tau,\mathcal{F}_{k}(\Theta),\|\cdot\|_{L_{2}(\mu)})\lesssim\log(1/\tau).$
Let $\tau=\varepsilon/4$, then we obtain that
$H_{B}(\varepsilon,\mathcal{F}_{k}(\Theta),\|.\|_{L^{2}(\mu)})\lesssim\log(1/\varepsilon).$
@@
$\mathcal{J}_{B}(\delta,\mathcal{F}_{k}(\Theta,\delta))=\int_{\delta^{2}/2^{13}}^{\delta}H_{B}^{1/2}(t,\mathcal{F}_{k}(\Theta,t),\|\cdot\|_{L_{2}(\mu)})~{}\mathrm{d}t\vee\delta\lesssim\int_{\delta^{2}/2^{13}}^{\delta}\log(1/t)dt\vee\delta.$ (32)
Let $\Psi(\delta)=\delta\cdot[\log(1/\delta)]^{1/2}$, then $\Psi(\delta)/\delta^{2}$ is a non-increasing function of $\delta$. Furthermore, equation (32) indicates that $\Psi(\delta)\geq\mathcal{J}_{B}(\delta,\mathcal{F}_{k}(\Theta,\delta))$. In addition, let $\delta_{n}=\sqrt{\log(n)/n}$, then we get that $\sqrt{n}\delta^{2}_{n}\geq c\Psi(\delta_{n})$ for some universal constant $c$. Finally, by applying Lemma D.1, we achieve the desired conclusion of the theorem.
@@
  • When $\alpha_{3}=e_{q,u}:=(0,\ldots,0,\underbrace{1}_{\textit{u-th}},0,\ldots,0)\in\mathbb{N}^{q}$ and $\ell_{1}=\mathbf{0}_{d}$, we have $\frac{1}{\mathcal{L}_{1n}}\cdot\sum_{i\in\mathcal{V}_{j}}\exp(c_{i}^{n})|(\Delta{\bm{\eta}}_{ij}^{n})^{(u)}|=|S_{n,j,\alpha_{3},\ell_{1}}|/\mathcal{L}_{1n}\to 0$ as $n\to\infty$. By taking the summation of the previous term with $u\in[q]$, we achieve that $\frac{1}{\mathcal{L}_{1n}}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i}^{n})\|\Delta{\bm{\eta}}_{ij}^{n}\|_{1}\to 0$. Owing to the topological equivalence between norm-1 and norm-2, it follows that

### v2

@@ Appendix D Proof of Theoretical Results
Here, the first inequality is according to the triangle inequality, the second inequality occurs as the softmax weight is bounded by 1, and the third inequality follows from the fact that the expert $h({\bm{x}},\cdot)$ is a Lipschitz function. Next, we have
$\|f_{\check{G}}-f_{\overline{G}}\|_{L^{\infty}}$ $=\Bigg{\|}\sum_{i=1}^{N}\Bigg{[}\frac{\exp({\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+({\bm{b}}_{i})^{\top}{\bm{x}}+c_{i})}{\sum_{j=1}^{k}\exp({\bm{x}}^{\top}{\bm{A}}_{j}{\bm{x}}+({\bm{b}}_{j})^{\top}{\bm{x}}+c_{j})}-\frac{\exp({\bm{x}}^{\top}\overline{{\bm{A}}}_{i}{\bm{x}}+(\overline{{\bm{b}}}_{i})^{\top}{\bm{x}}+\overline{c}_{i})}{\sum_{j=1}^{k}\exp({\bm{x}}^{\top}\overline{{\bm{A}}}_{j}{\bm{x}}+(\overline{{\bm{b}}}_{j})^{\top}{\bm{x}}+\overline{c}_{j})}\Bigg{]}\cdot h({\bm{x}},\overline{{\bm{\eta}}}_{i})\Bigg{\|}_{L^{\infty}}$
$\leq\sum_{i=1}^{N}\Bigg{\|}\Bigg{[}\frac{\exp({\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+({\bm{b}}_{i})^{\top}{\bm{x}}+c_{i})}{\sum_{j=1}^{k}\exp({\bm{x}}^{\top}{\bm{A}}_{j}{\bm{x}}+({\bm{b}}_{j})^{\top}{\bm{x}}+c_{j})}-\frac{\exp({\bm{x}}^{\top}\overline{{\bm{A}}}_{i}{\bm{x}}+(\overline{{\bm{b}}}_{i})^{\top}{\bm{x}}+\overline{c}_{i})}{\sum_{j=1}^{k}\exp({\bm{x}}^{\top}\overline{{\bm{A}}}_{j}{\bm{x}}+(\overline{{\bm{b}}}_{j})^{\top}{\bm{x}}+\overline{c}_{j})}\Bigg{]}\cdot h({\bm{x}},\overline{{\bm{\eta}}}_{i})\Bigg{\|}_{L^{\infty}}$
$\leq\sum_{i=1}^{N}\Bigg{\|}\frac{\exp({\bm{x}}^{\top}{\bm{A}}_{i}{\bm{x}}+({\bm{b}}_{i})^{\top}{\bm{x}}+c_{i})}{\sum_{j=1}^{k}\exp({\bm{x}}^{\top}{\bm{A}}_{j}{\bm{x}}+({\bm{b}}_{j})^{\top}{\bm{x}}+c_{j})}-\frac{\exp({\bm{x}}^{\top}\overline{{\bm{A}}}_{i}{\bm{x}}+(\overline{{\bm{b}}}_{i})^{\top}{\bm{x}}+\overline{c}_{i})}{\sum_{j=1}^{k}\exp({\bm{x}}^{\top}\overline{{\bm{A}}}_{j}{\bm{x}}+(\overline{{\bm{b}}}_{j})^{\top}{\bm{x}}+\overline{c}_{j})}\Bigg{\|}_{L^{\infty}}$
$\lesssim\sum_{i=1}^{N}\Big{[}\|{\bm{A}}_{i}-\overline{{\bm{A}}}_{i}\|\cdot\|{\bm{x}}\|^{2}+\|{\bm{b}}_{i}-\overline{{\bm{b}}}_{i}\|\cdot\|{\bm{x}}\|+|c_{i}-\overline{c}_{i}|\Big{]}$
$\leq\sum_{i=1}^{N}(\tau B^{2}+\tau B+\tau)\lesssim\tau.$
Above, the first inequality is due to the triangle inequality, the second inequality happens as the expert function is bounded, the third inequality follows from the fact that the softmax function is Lipschitz, and the fourth inequality occurs as the input space is bounded, that is, $\|{\bm{x}}\|\leq B$ for some constant $B>0$. According to the triangle inequality, we have
$\|f_{G}-f_{\overline{G}}\|_{L^{\infty}}\leq\|f_{G}-f_{\check{G}}\|_{L^{\infty}}+\|f_{\check{G}}-f_{\overline{G}}\|_{L^{\infty}}\lesssim\tau.$
@@
${N}(\tau,\mathcal{F}_{N}(\Theta),\|\cdot\|_{L^{\infty}})\leq|\Delta_{\tau}|\times|\Omega_{\tau}|\leq\mathcal{O}_{P}(n^{-(d^{2}+d+1)N})\times\mathcal{O}(n^{-qN})\leq\mathcal{O}(n^{-(d^{2}+d+1+q)N}).$ (31)
@@
$H_{B}(2\tau,\mathcal{F}_{N}(\Theta),\|\cdot\|_{L_{2}(\mu)})\lesssim\log(1/\tau).$
Let $\tau=\varepsilon/2$, then we obtain that
$H_{B}(\varepsilon,\mathcal{F}_{N}(\Theta),\|.\|_{L^{2}(\mu)})\lesssim\log(1/\varepsilon).$
@@
$\mathcal{J}_{B}(\delta,\mathcal{F}_{N}(\Theta,\delta))=\int_{\delta^{2}/2^{13}}^{\delta}H_{B}^{1/2}(t,\mathcal{F}_{N}(\Theta,t),\|\cdot\|_{L_{2}(\mu)})~{}\mathrm{d}t\vee\delta\lesssim\int_{\delta^{2}/2^{13}}^{\delta}\log(1/t)dt\vee\delta.$ (32)
Let $\Psi(\delta)=\delta\cdot[\log(1/\delta)]^{1/2}$, then $\Psi(\delta)/\delta^{2}$ is a non-increasing function of $\delta$. Furthermore, equation (32) indicates that $\Psi(\delta)\geq\mathcal{J}_{B}(\delta,\mathcal{F}_{N}(\Theta,\delta))$. In addition, let $\delta_{n}=\sqrt{\log(n)/n}$, then we get that $\sqrt{n}\delta^{2}_{n}\geq c\Psi(\delta_{n})$ for some universal constant $c$. Finally, by applying Lemma D.1, we achieve the desired conclusion of the theorem.
@@
  • When $\alpha_{3}=e_{q,u}:=(0,\ldots,0,\underbrace{1}_{\textit{u-th}},0,\ldots,0)\in\mathbb{N}^{q}$ and $\ell_{1}=\mathbf{0}_{d}$, we have
    $\frac{1}{\mathcal{L}_{1n}}\cdot\sum_{i\in\mathcal{V}_{j}}\exp(c_{i}^{n})|(\Delta{\bm{\eta}}_{ij}^{n})^{(u)}|=|S_{n,j,\alpha_{3},\ell_{1}}|/\mathcal{L}_{1n}\to 0\quad\text{as}\quad n\to\infty.$
    By taking the summation of the previous term with $u\in[q]$, we achieve that
    $\frac{1}{\mathcal{L}_{1n}}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i}^{n})\|\Delta{\bm{\eta}}_{ij}^{n}\|_{1}\to 0.$
    Owing to the topological equivalence between norm-1 and norm-2, it follows that